Find the equation of the line containing (4,-3)that is perpendicular to the line 6x-4y=1.
Answer by nerdybill(7384)

Find the equation of the line containing (4,-3)that is perpendicular to the line 6x-4y=1.
.
slope of:
6x-4y=1
-4y=-6x+1
4y=6x-1
y=(6/4)x-1/4
y=(3/2)x-1/4
slope is 3/2
.
A perpendicular line has a slope that is the negative reciprocal of the above:
new slope is -2/3
plug into "point-slope" form:
y - y1 = m(x - x1)
y - (-3) = (-2/3)(x - 4)
y + 3 = (-2/3)(x - 4)
y + 3 = (-2/3)x - (-2/3)4
y + 3 = (-2/3)x - (-8/3)
y + 3 = (-2/3)x + 8/3
y = (-2/3)x + 8/3 - 3
y = (-2/3)x + 8/3 - 9/3
y = (-2/3)x - 1/3